# RG theory of spontaneous stochasticity for Sabra model of turbulence

### v1

@@ I Introduction
Since the pioneering work of Lorenz [24] in 1969 it has been argued [23, 14, 8] that tiny errors, even as small as molecular noise [39], have a significant impact on the predictability of developed turbulent flows at moderate (turnover) times. Recent numerical studies suggested that although the deterministic description becomes unsuitable, the dynamics is still well defined as a stochastic process [29, 40, 3, 36]. The mathematical formulation of this stochastic process involves a limit in which the noise term in the equations of motion vanishes simultaneously with the viscous term. This limit yields a stochastic process that solves a formally deterministic initial value problem for an ideal flow described by the Euler equations. Such preservation of randomness in the zero-noise limit is called spontaneous stochasticity. A similar phenomenon was also detected in the description of particle trajectories carried by deterministic but rough (non-differentiable) velocity fields [5, 17, 21, 4]. Here the spontaneous stochasticity arises in the simultaneous limit of vanishing regularization of the velocity field and noise. In this limit trajectories remain stochastic despite they are carried by a formally deterministic velocity field. In both cases, a necessary (but not sufficient) condition is the non-uniqueness of solutions to the limiting (ideal) problem, and the stochastic description assumes a random choice of such non-unique solutions [16]. Spontaneous stochasticity of trajectories is often called Lagrangian, distinguishing it from Eulerian spontaneous stochasticity of velocity fields.
@@
Our goal here is to extend the RG approach to explain spontaneous stochasticity and its universality in shell turbulence models. Shell models are popular ”toy” models that share many non-trivial features of realistic turbulent flows [19, 6], and we focus on the so-called Sabra model [25]. Since the RG method was previously developed for discrete dynamics on a fractal lattice [33, 31], its extension to shell models faces two main obstacles: (a) the continuous-time formulation and (b) scale invariance with respect to both time and space scaling. Overcoming these difficulties is important not only within the framework of shell models, but also as a step towards the possible extension of the RG approach to realistic fluid systems. Developments in this direction for deterministic models (without noise) were presented in [32], but their extension to the stochastic setting is difficult. Instead, in the present paper we approximate the continuous-time model by dynamics on a properly designed spacetime lattice. This use of lattice models in situations where analysis in the continuous formulation is difficult or impossible is reminiscent of similar constructions in statistical mechanics and quantum field theory.
@@ II.1 Fluctuating Navier-Stokes equations
Here $\mathbf{u}(\mathbf{x},t)$ is a velocity field, which we consider in a triply $2\pi$-periodic domain $\mathbf{x}\in\mathbb{T}^{3}$, $p(\mathbf{x},t)$ is a pressure, and $\mathbf{f}(\mathbf{x},t)$ is a deterministic forcing applied at large scales. The fluctuating stress $\boldsymbol{\xi}$ is modeled as a Gaussian random matrix field with mean zero and covariance $\langle\xi_{ij}(\mathbf{x},t)\xi_{kl}(\mathbf{x}^{\prime},t^{\prime})\rangle=\left(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}-\frac{2}{3}\delta_{ij}\delta_{kl}\right)\delta^{3}(\mathbf{x}-\mathbf{x}^{\prime})\delta(t-t^{\prime})$. System (1) depends on two dimensionless parameters: $\mathrm{Re}$ and $\Theta$. Here $\mathrm{Re}=UL/\nu$ is the Reynolds number composed of the kinematic viscosity $\nu$ dividing the product of characteristic (large-scale) velocity $U$ and scale $L$. The second parameter is expressed in terms of thermodynamic properties of the fluid as $\Theta=2\nu k_{B}T/\rho L^{4}U^{3}$, where $k_{B}$ is Boltzmann’s constant, $T$ is absolute temperature, and $\rho$ is mass density. In addition, one can consider the Galerkin cutoff to Fourier modes with wave numbers $|\mathbf{k}|\leq 2\pi/\delta$, i.e. equate all modes with wave numbers $|\mathbf{k}|>2\pi/\delta$ to zero. Here $\delta=\ell_{\mathrm{mic}}/L$ refers to the microscopic (molecular) scale $\ell_{\mathrm{mic}}$ at which the macroscopic description (1) cease to be valid. Mathematically, the Galerkin truncation conviniently reduces Eq. (1) to a very large but still finite system of stochastic differential equations considered in the Fourier representation, thereby ensuring the well-posedness [2, 16]. Our focus will be on solving the initial value problem, where the initial condition
@@
The problem is that, as far as can be concluded from numerical studies, the turbulent solutions of the Navier-Stokes equations do not converge as $\mathrm{Re}\to\infty$.
@@
with the exponents $\beta=38/8$ and $\gamma=10/8$. This limit leads directly to the Euler equations (4), and we call any limit of this kind an ideal limit (generalizing the concept of an inviscid limit to include vanishing noise). The formulation (5) does not mean that the physical parameters change in a certain proportion. Rather, it is a proper mathematical approximation to the (very small, but finite) physical parameters (3).
@@
These four conditions are not formulated in precise mathematical form, since their explanation and validity for the three-dimensional fluctuating Navier-Stokes system are not well understood. These conditions imply that the stochastic process solving the fluctuating Navier-Stokes system converges in the ideal limit to a stochastic process solving the formally deterministic initial-value problem for the Euler system. A necessary requirement for this limiting process to exist is that the Euler system (4) and (2) is ill-posed, namely, has non-unique (weak) solutions [10, 9]. However, the existence of non-unique solutions does not in itself guarantee either convergence or the stochastic form of the limit. The third property is natural and can be justified mathematically [16]. The universality is a much stronger property: it means that the limiting stochastic process is precisely the same for a wide class of regularizations, types of noise and formulations of the ideal limit. For example, the limit should be independent of the exponents $\beta$ and $\gamma$ in Eq. (5), at least in a certain part of the parameter space.
Detailed and convincing evidence for spontaneous stochasticity is provided by shell models of turbulence, which are toy models of the Navier-Stokes system, as we discuss next. As for the fluctuating Navier-Stokes system itself, current evidence for spontaneous stochasticity is very limited. Convergence to a statistical solution in the ideal limit has been observed numerically for two-dimensional flows starting from the initial state of a vortex sheet, in close connection with the Kelvin-Helmholtz instability [18, 40]. In particular, in the work [40] the universality of the spontaneously stochastic solution was tested for two types of regularization: viscous and point vortex approximation. A recent study [36] confirms spontaneous stochasticity for the fluctuating Navier-Stokes system on a logarithmic lattice, which is another type of toy model.
@@ II.2 Fluctuating Sabra shell model
Note the last identity that follows from the expressions (14).
@@ II.4 Numerical observation of spontaneous stochasticity
Figure 1: (a) Evolution of the fluctuating Sabra model for $\mathrm{Re}=10^{9}$. Turbulent dynamics begins at the blowup time $t_{b}\approx 0.8$ (vertical line). (b) Turbulent state at the final time $T=1.5$.
@@ III Discrete-time fluctuating shell models
The Sabra shell model, although much simpler than the Navier-Stokes system, is still difficult to analyze numerically and theoretically. In this model, space is discretized, but time remains continuous and requires finite-difference approximation by numerical schemes. We now introduce a class of discrete-time models that, on the one hand, approximate the Sabra model to any accuracy (in the spirit of the finite-difference numerical method). On the other hand, such discrete-time models can themselves be considered as proper toy models of turbulence, since they preserve precisely the key properties: space-time scale invariance and energy conservation. These models have a considerable numerical advantage since their simulations are much faster and very accurate. Most importantly, they facilitate the explicit formulation of the renormalization group theory that explains the spontaneous stochasticity.
@@ III.1 Discrete-time algorithm
$u^{\prime}_{n}=u_{n}+B_{n}^{+}\left[\frac{\mathbf{u}+\mathbf{u}^{\prime}}{2}\right]\Delta\tau_{n},\quad u^{\prime}_{n+1}=u_{n+1}+B_{n+1}^{-}\left[\frac{\mathbf{u}+\mathbf{u}^{\prime}}{2}\right]\Delta\tau_{n},$ (22)
with all other components, $u^{\prime}_{m}=u_{m}$ for $m\neq n,n+1$, unchanged. Here the nonlinear terms (14) are evaluated for the mean state $(\mathbf{u}+\mathbf{u}^{\prime})/2$. The second relation in Eq. (22) is omitted at the cutoff shell $n=N$. Relations in Eq. (22) define the implicit finite-difference scheme for the nonlinear interaction in the Sabra model (6), in which the two parts of $B_{n}=B_{n}^{-}+B_{n}^{+}$ are assigned to different time steps. This scheme is designed in such a way that the energy is conserved exactly; see Appendix VIII.1, where Eqs. (22) are solved analytically. The dissipation and noise are introduced by solving the Ornstein-Uhlenbeck process
@@ III.2 Properties of discrete-time models
The specific design of our algorithm allows us to preserve several key properties of the Sabra model. First, the nonlinear scheme (22) preserves both the time and space scale invariance of the ideal system; see Eqs. (12) and (13). Second, it conserves the energy: $\mathcal{E}[\mathbf{u}]=\mathcal{E}[\mathbf{u}^{\prime}]$. Such precise conservation of energy is important for the stability of the algorithm, which otherwise may blowup in finite time. The energy changes at steps (25) as the work done by the boundary conditions, and dissipates due to viscosity at steps (24). Due to scale invariance and energy conservation, the discrete-time Sabra model features intermittency with anomalous scaling analogous to the Navier-Stokes turbulence [19]; see Appendix VIII.2. Thus, in addition to approximating the Sabra model, our discrete-time models are themselves full-fledged toy models of turbulence.
@@ III.3 Discrete-time LES shell model
$\frac{du_{n}}{dt}=-\nu_{n}u_{n}+\sigma_{n}\xi_{n},\quad\nu_{n}=Dk_{n}|u^{\prime}_{n}|,\quad\sigma_{n}=Ck_{n}^{1/2}|u^{\prime}_{n}|^{3/2},$ (26)
@@ IV RG relation for DT-LES models
In this section we study the dependence of the DT-LES models of the cutoff shell $N$. Our goal is to derive an explicit relation between solutions of this model for two successive cutoffs, $N$ and $N+1$. More specifically, we introduce the notion of a flow $\Phi^{(N)}$ that defines transition probabilities at any times and for any initial and boundary conditions. Then we derive the RG operator, $\mathcal{R}:\Phi^{(N)}\mapsto\Phi^{(N+1)}$, which expresses the flow $\Phi^{(N+1)}$ in terms of $\Phi^{(N)}$. This derivation is based on the discrete form and scaling symmetries of DT-LES models.
@@
The central idea is that the RG operator represents the ideal limit $N\to\infty$ as the RG dynamics in the space of flows, thereby linking limiting solutions to RG attractors. This theory has previously been developed for turbulence models on self-similar (fractal) lattices [33, 31], and here we extend it to more sophisticated shell models.
IV.1 Flow
@@
Let us consider the states at synchronized times $\boldsymbol{\tau}_{j}=(t_{j},\ldots,t_{j})$ with $j=0,1,2,\ldots$ and introduce the notation $\mathbf{s}_{j}=(\mathbf{u}_{j},t_{j})$; see Figs. 4 and 6. Then $\mathbf{s}_{0}=(\mathbf{u}_{\mathrm{ini}},0)$ is the initial state and the final state $\mathbf{s}_{\mathrm{fin}}=(\mathbf{u}_{\mathrm{fin}},T)$ is reached at some finite number of steps.
According to the algorithm of Section III.1, the transition between two subsequent states $\mathbf{s}_{j}$ and $\mathbf{s}_{j+1}$ can be expressed as a sequence of two maps. The first map is denoted as $\alpha_{T,\mathbf{b}}$ and defined by Eq. (25) with $\boldsymbol{\tau}=(t_{j},\ldots,t_{j})$ and $\Delta\tau_{1}$ given by Eq. (19). This map is deterministic and depends on the final time $T$ and boundary conditions $\mathbf{b}=(b_{-1},b_{0})$. The maps $\alpha_{T,\mathbf{b}}$ in the lattice dynamics are presented by loop arrows in Fig. 6. The second map is denoted as $\varphi_{T,\mathbf{b}}^{(N)}$ and defined by the composition of all steps (21), (22) and (24) leading to the next synchronized state with $\boldsymbol{\tau}=(t_{j+1},\ldots,t_{j+1})$. This map is random and depends on $T$, $\mathbf{b}$ and $N$. The maps $\varphi_{T,\mathbf{b}}^{(N)}$ in the lattice dynamics are presented by straight arrows in Fig. 6. Together, these two maps define the transition
$\varphi_{T,\mathbf{b}}^{(N)}\circ\alpha_{T,\mathbf{b}}:\mathbf{s}_{j}\mapsto\mathbf{s}_{j+1}.$ (27)
@@
Figure 6: Representation of the discrete-time algorithm of Fig. 4 in terms of the deterministic maps $\alpha_{T,\mathbf{b}}$ and the random maps $\varphi_{T,\mathbf{b}}^{(N)}$.
The random maps $\varphi_{T,\mathbf{b}}^{(N)}$ in Eq. (27) are statistically independent at each iteration. Hence, the sequence $\mathbf{s}_{0},\mathbf{s}_{1},\mathbf{s}_{2},\ldots$ represents a Markov chain, whose probability distribution we now describe. The transition probability $\mathbf{s}\mapsto\mathbf{s}_{\mathrm{new}}$ of a random map $\varphi_{T,\mathbf{b}}^{(N)}:\mathbf{s}\mapsto\mathbf{s}_{\mathrm{new}}$ is determined by a respective probability kernel $\Phi_{T,\mathbf{b}}^{(N)}(\mathbf{s}_{\mathrm{new}}|\mathbf{s})$. Mathematically, this kernel represents a probability distribution (probability measure) for $\mathbf{s}_{\mathrm{new}}$ given the initial state $\mathbf{s}$. It is also convenient to introduce the Dirac probability kernel $\mathrm{A}_{T,\mathbf{b}}(\mathbf{s}_{\mathrm{new}}|\mathbf{s})=\delta\big{(}\mathbf{s}_{\mathrm{new}}-\alpha_{T,\mathbf{b}}(\mathbf{s})\big{)}$ for the deterministic map $\alpha_{T,\mathbf{b}}:\mathbf{s}\mapsto\mathbf{s}_{\mathrm{new}}$. Then the transition probability of the composition (27) is given by the probability kernel $\big{(}\Phi_{T,\mathbf{b}}^{(N)}\circ\mathrm{A}_{T,\mathbf{b}}\big{)}(\mathbf{s}_{j+1}|\mathbf{s}_{j})$; see the Appendix VIII.3 for definitions of composition and other operations between probability kernels. The transition probability of $\mathbf{s}_{0}\mapsto\mathbf{s}_{m}$, which includes $m$ steps of the Markov chain, is determined by the probability kernel
@@
The number of steps $m$ between the initial state $\mathbf{s}_{0}=(\mathbf{u}_{\mathrm{ini}},0)$ and the final state $\mathbf{s}_{\mathrm{fin}}=(\mathbf{u}_{\mathrm{fin}},T)$ is a random number. Note that $\mathbf{s}_{\mathrm{fin}}$ is the stationary state of the Markov chain (27), because the algorithm is designed to stop at a given time $T$. The transition probability $\mathbf{u}_{\mathrm{ini}}\mapsto\mathbf{u}_{\mathrm{fin}}$ is determined by the respective probability kernel, which we denote as $\mathrm{P}_{T,\mathbf{b}}^{(N)}(\mathbf{u}_{\mathrm{fin}}|\mathbf{u}_{\mathrm{ini}})$. Using Eq. (28) we express
$\mathrm{P}_{T,\mathbf{b}}^{(N)}(\mathbf{u}_{\mathrm{fin}}|\mathbf{u}_{\mathrm{ini}})=\big{(}\Phi_{T,\mathbf{b}}^{(N)}\circ\mathrm{A}_{T,\mathbf{b}}\big{)}^{\infty}(\mathbf{s}_{\mathrm{fin}}|\mathbf{s}_{0}),$ (29)
where the sequence of compositions converges in a finite (but random) number of steps.
We will use the short notation
$\Phi^{(N)}=\{\Phi_{T,\mathbf{b}}^{(N)}:T\geq 0,\,\mathbf{b}\in\mathbb{C}^{2}\}$ (30)
for the family of kernels with arbitrary times and boundary conditions. This family defines all transition probabilities in the DT-LES model using Eq. (29). Denoting the family of these transition probabilities as $\mathrm{P}^{(N)}=\{\mathrm{P}_{T,\mathbf{b}}^{(N)}:T\geq 0,\,\mathbf{b}\in\mathbb{C}^{2}\}$, we write
@@
where the operator $\mathcal{P}$ denotes the transformation (29). By analogy with the theory of dynamical systems, we call $\Phi^{(N)}$ a flow. We conclude that the flow $\Phi^{(N)}$ defines all transition probabilities $\mathrm{P}^{(N)}$ in the DT-LES model for a given cutoff $N$.
IV.2 RG relation between the flows
In this section, we derive the RG operator, which expresses the flow $\Phi^{(N+1)}$ in terms of $\Phi^{(N)}$. Since we are dealing with different $N$, it is convenient to think of the states $\mathbf{u}=(u_{1},u_{2},\ldots)$ as infinite sequences, which are filled by zeros at the shells $n>N$. Let us first introduce some deterministic maps that are needed to formulate our main result:

### v2

@@ I Introduction
Since the pioneering work of Lorenz [27] in 1969 it has been argued [25, 17, 8, 26] that tiny errors, even as small as molecular noise [42], have a significant impact on the predictability of developed turbulent flows at moderate (turnover) times. Recent numerical studies suggested that although the deterministic description becomes unsuitable, the dynamics is still well defined as a stochastic process [33, 43, 3, 39]. The mathematical formulation of this stochastic process involves a limit in which the noise term in the equations of motion vanishes simultaneously with the viscous term. This limit yields a stochastic process that solves a formally deterministic initial value problem for an ideal flow described by the Euler equations. Such preservation of randomness in the zero-noise limit is called spontaneous stochasticity. A similar phenomenon was also detected in the description of particle trajectories carried by deterministic but rough (non-differentiable) velocity fields [5, 18, 22, 4]. Here the spontaneous stochasticity arises in the simultaneous limit of vanishing regularization of the velocity field and noise. In this limit trajectories remain stochastic despite they are carried by a formally deterministic velocity field. In both cases, a necessary (but not sufficient) condition is the non-uniqueness of solutions to the limiting (ideal) problem, and the stochastic description assumes a random choice of such non-unique solutions [16]. Spontaneous stochasticity of trajectories is often called Lagrangian, distinguishing it from Eulerian spontaneous stochasticity of velocity fields.
@@
Our goal here is to extend the RG approach to explain spontaneous stochasticity and its universality in shell turbulence models. Shell models are popular ”toy” models that share many non-trivial features of realistic turbulent flows [20, 7], and we focus on the so-called Sabra model [23]. Since the RG method was previously developed for discrete-time dynamics on a fractal lattice [28, 35], its extension to shell models requires addressing two key issues: the continuous-time formulation and scale invariance with respect to both time and space. This extension is important not only within the framework of shell models, but also as a step toward a possible application of the RG approach to more realistic fluid systems. Developments in this direction for deterministic models (without noise) were presented in [36]; however, a similar extension to the stochastic setting faces significant mathematical difficulties in defining the RG operator. Instead, in the present paper we approximate the continuous-time model by dynamics on a properly designed spacetime lattice. This use of lattice models in situations where analysis in the continuous formulation is problematic is reminiscent of analogous constructions in statistical mechanics and quantum field theory.
@@ II.1 Fluctuating Navier-Stokes equations
Here $\mathbf{u}(\mathbf{x},t)$ is a velocity field, which we consider in a triply $2\pi$-periodic domain $\mathbf{x}\in\mathbb{T}^{3}$, $p(\mathbf{x},t)$ is a pressure, and $\mathbf{f}(\mathbf{x},t)$ is a deterministic forcing applied at large scales. The fluctuating stress $\boldsymbol{\xi}$ is modeled as a Gaussian random matrix field with mean zero and covariance $\langle\xi_{ij}(\mathbf{x},t)\xi_{kl}(\mathbf{x}^{\prime},t^{\prime})\rangle=\left(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}-\frac{2}{3}\delta_{ij}\delta_{kl}\right)\delta^{3}(\mathbf{x}-\mathbf{x}^{\prime})\delta(t-t^{\prime})$. System (1) depends on two dimensionless parameters: $\mathrm{Re}$ and $\Theta$. Here $\mathrm{Re}=UL/\nu$ is the Reynolds number composed of the kinematic viscosity $\nu$ dividing the product of characteristic (large-scale) velocity $U$ and scale $L$. The second parameter is expressed in terms of thermodynamic properties of the fluid as $\Theta=2\nu k_{B}T/\rho L^{4}U^{3}$, where $k_{B}$ is Boltzmann’s constant, $T$ is absolute temperature, and $\rho$ is mass density. In addition, one can consider the Galerkin cutoff to Fourier modes with wave numbers $|\mathbf{k}|\leq 2\pi/\delta$, i.e. equate all modes with wave numbers $|\mathbf{k}|>2\pi/\delta$ to zero. Here $\delta=\ell_{\mathrm{mic}}/L$ refers to the microscopic (molecular) scale $\ell_{\mathrm{mic}}$ at which the macroscopic description (1) cease to be valid. Mathematically, the Galerkin truncation conveniently reduces Eq. (1) to a very large but still finite system of stochastic differential equations considered in the Fourier representation, thereby ensuring the well-posedness [2, 16]. In particular, the gradient of the noise term in Eq. (1) is well defined in this truncated Fourier representation.
Our focus will be on solving the initial value problem, where the initial condition
@@
The problem is that, based on numerical evidence, turbulent solutions of the Navier–Stokes equations do not appear to converge as $\mathrm{Re}\to\infty$.
@@
with the exponents $\beta=38/8$ and $\gamma=10/8$. This limit leads directly to the Euler equations (4), and we refer to any limit of this type as an ideal limit, generalizing the concept of the inviscid limit to include vanishing noise. The formulation (5) does not mean that the physical parameters vary in a fixed proportion. Rather, it provides a proper mathematical approximation to the (very small, but finite) physical parameters (3).
@@
These conditions imply that the stochastic process solving the fluctuating Navier-Stokes system converges, in the ideal limit, to a stochastic process that solves the formally deterministic initial-value problem for the Euler equations. A necessary requirement for this limiting process to exist is that the Euler system (4) and (2) is ill-posed, namely, has non-unique (weak) solutions [11, 9]. However, the mere existence of non-unique solutions does not, by itself, guarantee either convergence or a stochastic form of the limit. The third property is natural and can be justified mathematically [16]. Universality is a much stronger property: it means that the limiting stochastic process is identical for a wide class of regularizations, noise types, and formulations of the ideal limit. For example, the limit should be independent of the exponents $\beta$ and $\gamma$ in Eq. (5), at least within certain intervals of their values. The reader should regard our characterization of spontaneous stochasticity more as a guideline than as a proposal for a precise mathematical formulation. Indeed, its interpretation and validity for the three-dimensional fluctuating Navier-Stokes system remain poorly understood, while simplified models suggest a broader range of possible scenarios [35].
Detailed and convincing evidence for spontaneous stochasticity is provided by shell models of turbulence, which are toy models of the Navier-Stokes system, as we discuss next. As for the fluctuating Navier-Stokes system itself, convergence to a statistical solution in the ideal limit has been observed numerically for two-dimensional flows starting from the initial state of a vortex sheet, in close connection with the Kelvin-Helmholtz instability [19, 43]. In particular, in the work [43] the universality of the spontaneously stochastic solution was tested for two types of regularization: viscous and point vortex approximation. A recent study [39] confirms spontaneous stochasticity for the fluctuating Navier-Stokes system on a logarithmic lattice, which is another type of toy model.
@@ II.2 Fluctuating Sabra shell model
Note the last identity, which follows from the expressions in Eq. (14).
@@ II.4 Numerical observation of spontaneous stochasticity
Figure 1: (a) Evolution of the fluctuating Sabra model for $\mathrm{Re}=10^{9}$. Turbulent dynamics begins at the blowup time $t_{b}\approx 0.8$ (vertical line). Each curve corresponds to a different shell $n=1,\ldots,N$ with $N=35$. (b) Turbulent state at the final time $T=1.5$ for the realization from the left panel.
@@ III Discrete-time fluctuating shell models
The Sabra shell model, although much simpler than the Navier-Stokes system, is still difficult to analyze numerically and theoretically. In this model, space is discretized, but time remains continuous and requires finite-difference approximation by numerical schemes. We now introduce a class of discrete-time models that, on the one hand, approximate the Sabra model to any accuracy (in the spirit of the finite-difference numerical method). On the other hand, such discrete-time models can themselves be considered as proper toy models of turbulence, since they preserve precisely the key properties: space-time scale invariance and energy conservation. These models have a considerable numerical advantage since their simulations are much faster and very accurate. Most importantly, their conservative and symmetry-preserving structure, which is not retained by standard numerical schemes, allows for an explicit and transparent formulation of the renormalization-group theory.
@@ III.1 Discrete-time algorithm
$u^{\prime}_{n}=u_{n}+B_{n}^{+}\left[\frac{\mathbf{u}+\mathbf{u}^{\prime}}{2}\right]\Delta\tau_{n},\quad u^{\prime}_{n+1}=u_{n+1}+B_{n+1}^{-}\left[\frac{\mathbf{u}+\mathbf{u}^{\prime}}{2}\right]\Delta\tau_{n}.$ (22)
Note that this relation modifies only the two components at shells $n$ and $n+1$, while all other components are left unchanged, so that $u^{\prime}_{m}=u_{m}$ for $m\neq n,n+1$. Here the nonlinear terms (14) are evaluated for the mean state $(\mathbf{u}+\mathbf{u}^{\prime})/2$. The second relation in Eq. (22) is omitted at the cutoff shell $n=N$. Relations in Eq. (22) define the implicit finite-difference scheme for the nonlinear interaction in the Sabra model (6), in which the two parts $B_{n}^{-}$ and $B_{n}^{+}$ of the nonlinear term are treated independently. This scheme is designed in such a way that the energy is conserved exactly; see Appendix VIII.1, where Eqs. (22) are solved analytically. The dissipation and noise are introduced by solving the Ornstein-Uhlenbeck process
@@
We refer the reader to Appendix VIII.2 for an explicit end-to-end description of the algorithm, ensuring its reproducibility on a computer. There we also justify that the algorithm converges in a finite—though random—number of steps.
@@ III.2 Properties of discrete-time models
The specific design of our algorithm allows us to preserve several key properties of the Sabra model. First, the nonlinear scheme (22) preserves both the time and space scale invariance of the ideal system; see Eqs. (12) and (13). Second, it conserves the energy: $\mathcal{E}[\mathbf{u}]=\mathcal{E}[\mathbf{u}^{\prime}]$. Such precise conservation of energy is important for the stability of the algorithm, which otherwise may blowup in finite time. The energy changes at steps (25) as the work done by the boundary conditions, and dissipates due to viscosity at steps (24). Due to scale invariance and energy conservation, the discrete-time Sabra model features intermittency with anomalous scaling analogous to the Navier-Stokes turbulence [20]; see Appendix VIII.3. Thus, in addition to approximating the Sabra model, our discrete-time models are themselves full-fledged toy models of turbulence.
@@ III.3 Discrete-time LES shell model
$\frac{du_{n}}{dt}=-\nu_{n}u_{n}+\sigma_{n}\eta_{n},\quad\nu_{n}=Dk_{n}|u^{\prime}_{n}|,\quad\sigma_{n}=Ck_{n}^{1/2}|u^{\prime}_{n}|^{3/2},$ (26)
@@ IV RG relation for DT-LES models
In this section we study the dependence of the DT-LES models of the cutoff shell $N$. Our goal is to derive an explicit relation between solutions of this model for two successive cutoffs, $N$ and $N+1$. More specifically, we introduce the notion of a flow kernel $\Phi^{(N)}$ that defines transition probabilities at any times and for any initial and boundary conditions. Then we derive the RG operator, $\mathcal{R}:\Phi^{(N)}\mapsto\Phi^{(N+1)}$, which expresses the flow kernel $\Phi^{(N+1)}$ in terms of $\Phi^{(N)}$. This derivation is based on the discrete form and scaling symmetries of DT-LES models.
@@
The central idea is that the RG operator represents the ideal limit $N\to\infty$ as the RG dynamics in the space of flow kernels, thereby linking limiting solutions to RG attractors. This theory has previously been developed for turbulence models on self-similar (fractal) lattices [28, 35], and here we extend it to more sophisticated shell models.
IV.1 Flow kernel
@@
Let us consider the states at synchronized times $\boldsymbol{\tau}_{j}=(t_{j},\ldots,t_{j})$, where the index $j=0,1,2,\ldots$ labels the time steps at the shell $n=1$; see Figs. 4 and 6. We introduce the notation $\mathbf{s}_{j}=(\mathbf{u}_{j},t_{j})$, which collects the values of the state variables and the corresponding time into a single vector. Then $\mathbf{s}_{0}=(\mathbf{u}_{\mathrm{ini}},0)$ is the initial state and we denote the final state by $\mathbf{s}_{\mathrm{fin}}=(\mathbf{u}_{\mathrm{fin}},T)$.
[…]
$\big(\Phi_{T,\mathbf{b}}^{(N)}\circ\mathrm{A}_{T,\mathbf{b}}\big)(\mathbf{s}_{j+1}|\mathbf{s}_{j}),$ (27)
which is the same for all values of $j$. Here $\mathrm{A}_{T,\mathbf{b}}(\tilde{\mathbf{s}}|\mathbf{s}_{j})=\delta\big(\tilde{\mathbf{s}}-\alpha_{T,\mathbf{b}}(\mathbf{s}_{j})\big)$ is the Dirac kernel corresponding to the map $\alpha_{T,\mathbf{b}}$, and we refer to Appendix VIII.4 for definitions of composition and other operations between probability kernels.
@@
Figure 6: Representation of the discrete-time algorithm of Fig. 4 in terms of deterministic maps $\alpha_{T,\mathbf{b}}$ and stochastic transitions governed by the flow kernels $\Phi_{T,\mathbf{b}}^{(N)}$.
Observe that the transitions $\mathbf{s}_{j}\mapsto\mathbf{s}_{j+1}$ for different values of $j$ are statistically independent. Hence, the transition probability from $\mathbf{s}_{0}$ to $\mathbf{s}_{m}$, which consists of $m$ independent steps, is determined by the probability kernel
@@
The number of steps $m$ between the initial state $\mathbf{s}_{0}=(\mathbf{u}_{\mathrm{ini}},0)$ and the final state $\mathbf{s}_{\mathrm{fin}}=(\mathbf{u}_{\mathrm{fin}},T)$ is a random variable. Treating the attainment of the final state as a stopping condition, we denote the corresponding kernel by
$\big(\Phi_{T,\mathbf{b}}^{(N)}\circ\mathrm{A}_{T,\mathbf{b}}\big)^{\mathrm{fin}},$ (29)
with the understanding that a sufficient number of iterations is performed to reach the final time $T$. Dropping the temporal components of the states $\mathbf{s}_{0}$ and $\mathbf{s}_{\mathrm{fin}}$, which are now redundant, we obtain the transition probability for $\mathbf{u}_{\mathrm{ini}}\mapsto\mathbf{u}_{\mathrm{fin}}$ in the form of a Markov kernel
$\mathrm{P}_{T,\mathbf{b}}^{(N)}(\mathbf{u}_{\mathrm{fin}}|\mathbf{u}_{\mathrm{ini}})=\big(\Phi_{T,\mathbf{b}}^{(N)}\circ\mathrm{A}_{T,\mathbf{b}}\big)^{\mathrm{fin}}(\mathbf{s}_{\mathrm{fin}}|\mathbf{s}_{0}).$ (30)
We will use the short notations
$\Phi^{(N)}=\{\Phi_{T,\mathbf{b}}^{(N)}:T\geq 0,\,\mathbf{b}\in\mathbb{C}^{2}\},\quad\mathrm{P}^{(N)}=\{\mathrm{P}_{T,\mathbf{b}}^{(N)}:T\geq 0,\,\mathbf{b}\in\mathbb{C}^{2}\}$ (31)
for the families of kernels with arbitrary times and boundary conditions. They are related as
@@
where the operator $\mathcal{P}$ denotes the transformation (30). By analogy with the theory of dynamical systems, we call $\Phi^{(N)}$ a flow kernel. We conclude that the flow kernel $\Phi^{(N)}$ defines all transition probabilities $\mathrm{P}^{(N)}$ in the DT-LES model for a given cutoff $N$.
IV.2 RG relation between the flow kernels
In this section, we derive the RG operator, which expresses the flow kernel $\Phi^{(N+1)}$ in terms of $\Phi^{(N)}$. Since we are dealing with different $N$, it is convenient to think of the states $\mathbf{u}=(u_{1},u_{2},\ldots)$ as infinite sequences, which are filled by zeros at the shells $n>N$. Let us first introduce some deterministic maps that are needed to formulate our main result: